lisa has 5 skirts 6 blouses and 3 jackets how many three piece outfits can she put together assuming any piece can go with any other?
Answer by jim_thompson5910(35256)

Multiply out all the numbers to get 5*6*3 = 30*3 = 90 different possible outfits.